For a certain ideal gas Cp = 1.1 KJ/kg-K and k = 1.3 , Determine:
a. Molecular weight
b. Universal gas constant
c. Constant volume specific heat
1
Expert's answer
2021-10-07T08:49:11-0400

c)

"k=\frac{c_p}{c_V},\implies" "c_V=\frac{c_p}k=0.27~\frac{kJ}{kg\cdot K},"

b)

"R=c_p-c_V=8.3~\frac J{kg\cdot K},"

a)

"n=\frac{c_p-c_V}R=10."